A father is 41 years old and his son is 9. In how many years will the father be three times as old as his son?
Answer by tanjo3(60)

Let n be the number of years until the stated condition is met, then:

41 + n = 3(9 + n)
41 + n = 27 + 3n
14 = 2n
7 = n

In 7 years, the father will be three times as old as his son (he will be 48 and his son will be 16).